Compare the amount of heat given off by 1.40 mole of liquid water when it cools from 100.0 oC to 30.0 oC to

    that given off when 1.40 mol of steam cools from 200.0 oC to 110.0 oC   ( Cp H2O (l)  = 4.184 j/g-oC  , Cp H2O (g)

    = 1.87 J/j-oC) . Explain your comparison


1
Expert's answer
2022-04-21T13:43:03-0400

Mass of water = 1.40 * 18 = 25.2 (g)

Q1 = cmΔT = 4.184 * 25.2 * (100.0 – 30.0) = 7380.6 J

Q2 = 1.87 * 25.2 * (200.0 – 110.0) = 4241.2 J

Q1 > Q2